OpenStudy (fanduekisses):

I'm having trouble with this equation: (x-5)/2 + (1/2)=2x-(x-3)/8

OpenStudy (anonymous):

what's the trouble? ^.^ denominators get you down?

OpenStudy (fanduekisses):

I' confused lol I need a little refresher

OpenStudy (anonymous):

\[\Large \frac{x-5}2+\frac12= 2x - \frac{x-3}{8}\]

OpenStudy (fanduekisses):

What do I do? Do I have to find coon denominator?

OpenStudy (anonymous):

Yeah, I guess ^.^ You could multiply everything by the LCD

OpenStudy (fanduekisses):

oh, which is 2 right?

OpenStudy (anonymous):

No... ^.^

OpenStudy (fanduekisses):

D:

OpenStudy (fanduekisses):

please help

OpenStudy (anonymous):

It's that number, which, if you multiply everything with it, all the denominators just... disappear ^.^

OpenStudy (fanduekisses):

0?

OpenStudy (anonymous):

LOL that actually works, but that doesn't help.... What are all the denominators?

OpenStudy (fanduekisses):

2 2's and one 8

OpenStudy (anonymous):

2 and 8... now what's their least common multiple?

OpenStudy (fanduekisses):

8

OpenStudy (anonymous):

That's right ^.^ Now multiply everything (in your equation) by 8.

OpenStudy (fanduekisses):

8x-40+8=16x-8x-24?

OpenStudy (anonymous):

This'll take a while >.>

OpenStudy (anonymous):

Never heard of cancellation?

OpenStudy (fanduekisses):

yes

OpenStudy (anonymous):

Well then, redo it :)

OpenStudy (fanduekisses):

when multiplying 8 times \[\frac{ x-5}{ 2 }\] do i multiply the x too?

OpenStudy (anonymous):

8 cancels out the 2, and leaves...?

OpenStudy (fanduekisses):

x-5

OpenStudy (fanduekisses):

:}

OpenStudy (anonymous):

No... come on :( What's 8 divided by 2?

OpenStudy (fanduekisses):

4

OpenStudy (fanduekisses):

so, \[\frac{ x-5 }{4 }\] ?